# Vision-Based Mobile App GUI Testing: A Survey

### v2

@@ Abstract.
  Graphical User Interface (GUI) has become one of the most significant parts of mobile applications (apps). It is a direct bridge between mobile apps and end users, which directly affects the end user’s experience. Neglecting GUI quality can undermine the value and effectiveness of the entire mobile app solution. Significant research efforts have been devoted to GUI testing, one effective method to ensure mobile app quality. By conducting rigorous GUI testing, developers can ensure that the visual and interactive elements of the mobile apps not only meet functional requirements but also provide a seamless and user-friendly experience. However, traditional solutions, relying on the source code or layout files, have met challenges in both effectiveness and efficiency due to the gap between what is obtained and what app GUI actually presents. Vision-based mobile app GUI testing approaches emerged with the development of computer vision technologies and have achieved promising progress. In this survey paper, we provide a comprehensive investigation of the state-of-the-art techniques on 271 papers, among which 92 are vision-based studies. This survey covers different topics of GUI testing, like GUI test generation, GUI test record & replay, GUI testing framework, etc. Specifically, the research emphasis of this survey is placed mostly on how vision-based techniques outperform traditional solutions and have gradually taken a vital place in the GUI testing field. Based on the investigation of existing studies, we outline the challenges and opportunities of (vision-based) mobile app GUI testing and propose promising research directions with the combination of emerging techniques.
@@
copyright: acmcopyrightjournalyear: 2023doi: 1journal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Mobile app testing is a widely investigated topic. Some studies on mobile app GUI testing focus on automated GUI testing or analysis of the overall development path of GUI testing (Arnatovich and Wang, 2018; Banerjee et al., 2013; Said et al., 2020). However, such surveys organize the literature according to the paper topic, ignoring the important approach basis for retrieving app state information, which is a fundamental factor for mobile app GUI testing. Our survey focuses on the source code requirement (i.e., black-box, grey-box, and white-box) and approach basis (i.e., code, layout, and image) to explore the development of mobile app GUI testing techniques over years.
There have been also more and more studies investigating the application of advanced technologies in different software engineering tasks. Bajammal et al. (Bajammal et al., 2020) explores the application of CV in general areas of software engineering. Some other studies focus on discovering the application of new techniques (e.g., deep learning) on general software (Linares-Vásquez et al., 2017; Ricca et al., 2021). The survey presented by Wang et al. in 2024 (Wang et al., 2024) is a recent and representative work that primarily focuses on the application of large language models (LLMs) in software testing, providing a comprehensive review of the use of LLMs in various software testing tasks, such as unit test generation, program repair, and fuzz testing. However, mobile app GUI testing is not well emphasized, investigated, or discussed. To supplement the missing investigation on how mobile app GUI testing develops with the emerging of advanced techniques, and how it develops from non-vision-based techniques to vision-based ones, our survey investigates the overall development of mobile app GUI testing, spanning eight major topics in mobile app GUI testing for in-depth research, with a special focus on the approach basis, including non-vision-based and vision-based ones. Most importantly, we are the first review to focus on research on vision-based mobile app GUI testing techniques.
Table 1. Mobile App GUI Testing Research Topics
Topic Abbreviation # Total Paper # Vision-based Paper
GUI Test Generation GEN 62 10
GUI Test Record & Replay (Migration) R&R 34 13
GUI Testing Framework FRM 56 8
GUI Test Script Maintenance SRP 16 3
GUI Test Report Analysis RPT 34 14
GUI Element Detection ELE 39 36
GUI Usability & Accessibility Testing U&A 11 6
GUI Testing Evaluation Criteria EVA 19 2
Total 271 92
@@
We believe that this survey presents a unique research perspective in the field of vision-based mobile app GUI testing. We, for the first time, focus on vision-based mobile app GUI testing techniques, which have been relatively under-explored in previous secondary research; thus, we consider this in itself to be a novel contribution. Through this unique perspective, we hope to fill the research gap in this field in current reviews and provide a more comprehensive and in-depth understanding of the development of vision-based GUI testing. In this survey, we not only introduce the basic concepts and methods of vision-based mobile app GUI testing but also delve into several key topics within this field, including GUI test generation, GUI test record & replay, GUI testing framework, GUI test script maintenance, GUI test report analysis, GUI element detection, GUI usability and accessibility testing, and GUI testing evaluation criteria. The exploration of these topics is based on an in-depth analysis and summary of relevant academic literature, aiming to provide readers with a comprehensive and profound understanding. Furthermore, we highlight the uniqueness and advantages of vision-based GUI testing by comparing it with other non-vision-based GUI testing methods. Therefore, we believe that the paper has made substantial contributions to the novelty of vision-based mobile app GUI testing.
We present existing studies on mobile app GUI testing from the following eight topics, as illustrated in Table 1  [^1] We use the abbreviations in the rest of this paper to denote the topics.:
@@
GUI Test Record & Replay (Migration) (R&R): This topic targets at maximizing the value of manually written test scripts, which enables the automated execution of test scripts on different running environments without manual adaptions. Such test scripts contain the domain knowledge of human testers and can better test the critical functionalities. It also addresses the challenge of ensuring compatibility across different platforms and devices, facilitating testing in diverse environments.

### v3

@@ Abstract.
Graphical User Interface (GUI) has become one of the most significant parts of mobile applications (apps). It is a direct bridge between mobile apps and end users, which directly affects the end user’s experience. Neglecting GUI quality can undermine the value and effectiveness of the entire mobile app solution. Significant research efforts have been devoted to GUI testing, one effective method to ensure mobile app quality. By conducting rigorous GUI testing, developers can ensure that the visual and interactive elements of the mobile apps not only meet functional requirements but also provide a seamless and user-friendly experience. However, traditional solutions, relying on the source code or layout files, have met challenges in both effectiveness and efficiency due to the gap between what is obtained and what app GUI actually presents. Vision-based mobile app GUI testing approaches emerged with the development of computer vision technologies and have achieved promising progress. In this survey paper, we provide a comprehensive investigation of the state-of-the-art techniques on 271 papers, among which 92 are vision-based studies. This survey covers different topics of GUI testing, like GUI test generation, GUI test record & replay, GUI testing framework, etc. In particular, we highlight the emerging role of vision-based techniques and analyze how they reshape traditional approaches to mobile app GUI testing. Based on the investigation of existing studies, we outline the challenges and opportunities of (vision-based) mobile app GUI testing and propose promising research directions with the combination of emerging techniques.
@@
copyright: acmcopyrightjournalyear: 2025doi: 1journal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Mobile app testing is a widely studied topic. Several surveys on mobile app GUI testing emphasize automated testing or the overall evolution of GUI testing (Arnatovich and Wang, 2018; Banerjee et al., 2013; Said et al., 2020). Several of related studies offer valuable perspectives. For example, a bibliometric overview that charts the growth of GUI testing over time (Rodríguez-Valdés et al., 2021), the survey focused on Android app testing techniques (Li, 2023), or reviews centered around user experience and the human-centric aspects of mobile GUI evaluation (Deshmukh et al., 2023). Our survey addresses this gap by focusing on source code requirements (i.e., black-box, grey-box, white-box) and approach basis (i.e., code, layout, image) to trace the evolution of mobile app GUI testing techniques over time. While these surveys offer valuable insights, they differ from our work in scope and focus. Most target general GUI testing, Android-specific issues, or UX considerations, without addressing the broader evolution of mobile GUI testing. Crucially, none examine the emerging role of vision-based techniques or systematically compare them with traditional approaches across core dimensions. Our study is the first, to our knowledge, to provide a comparative and cross-cutting perspective that highlights this shift toward visual reasoning and its implications for future research and tool development. Recent studies have also explored the integration of advanced technologies in software engineering. For instance, Bajammal et al. (Bajammal et al., 2020) examine the application of computer vision (CV) in general software engineering, while others investigate the use of emerging techniques (e.g., deep learning) in software development (Linares-Vásquez et al., 2017; Ricca et al., 2021). Wang et al. (Wang et al., 2024) provide a comprehensive overview of large language models (LLMs) in software testing. However, these surveys lack emphasis on mobile app GUI testing. To address this, our work provides a detailed investigation into the development of mobile app GUI testing, highlighting the transition from non-vision-based to vision-based techniques. Covering eight major topics, we categorize the literature by approach basis and source code requirement. Notably, we present the first comprehensive review focused on vision-based mobile app GUI testing techniques.
Table 1. Mobile App GUI Testing Research Topics and Approach Taxonomy
Topic # Total Paper # Vision-based Paper Source Code Requirement Approach Basis (C: Code, L: Layout, I: Image)
Black-box Grey-box White-box C L I C & L L & I C & L & I
GEN 62 10 28 21 13 32 18 10 2 0 0
R&R 34 13 20 11 3 12 8 12 1 1 0
FRM 56 8 33 14 9 24 24 8 0 0 0
SRP 16 3 6 5 5 7 4 2 2 1 0
RPT 34 14 30 1 3 17 3 14 0 0 0
ELE 39 36 32 6 1 2 0 31 0 5 1
U&A 11 6 9 2 0 1 3 5 1 1 0
EVA 19 2 12 1 6 7 11 1 0 0 0
Sum 271 92 170 61 40 102 71 83 6 8 1
@@
We believe that this survey presents a unique research perspective in the field of vision-based mobile app GUI testing. We, for the first time, focus on vision-based mobile app GUI testing techniques, which have been relatively under-explored in previous secondary research; thus, we consider this in itself to be a novel contribution. Through this unique perspective, we hope to fill the research gap in this field in current reviews and provide a more comprehensive and in-depth understanding of the development of vision-based GUI testing. In this study, we aim to systematically review the landscape of mobile app GUI testing techniques, with a particular emphasis on the emergence of vision-based approaches. Vision-based testing leverages visual elements, such as screenshots and rendered UI states, rather than relying solely on structural or code-based representations. Our goal is to understand how the field of mobile GUI testing has evolved, and how vision-based methods differ from traditional techniques in addressing long-standing problems. To this end, we identify and classify techniques across various testing dimensions, highlighting trends and future directions driven by visual reasoning. While our analysis covers the general landscape of mobile GUI testing, vision-based techniques are not discussed in isolation but are integrated into each aspect of the review. In this survey, we not only introduce the basic concepts and methods of vision-based mobile app GUI testing but also delve into several key topics within this field (see Table 1  [^1] The full topic names are shown below. We use the abbreviations in the rest of this paper to denote the topics.), including GUI test generation, GUI test record & replay, GUI testing framework, GUI test script maintenance, GUI test report analysis, GUI element detection, GUI usability and accessibility testing, and GUI testing evaluation criteria. The exploration of these topics is based on an in-depth analysis and summary of relevant academic literature, aiming to provide readers with a comprehensive and profound understanding. Furthermore, we highlight the uniqueness and advantages of vision-based GUI testing by comparing it with other non-vision-based methods. We highlight the emerging role of vision-based techniques and analyze how they reshape traditional approaches to mobile app GUI testing. Therefore, we believe that the paper has made substantial contributions to the novelty of vision-based mobile app GUI testing.
In constructing our topic taxonomy, we adopt a hybrid top-down and bottom-up approach. From a top-down perspective, we refer to SWEBOK V4.0, Section 6.2: Testing in the Application Domains, which explicitly discusses GUI testing as a domain requiring specialized techniques and tools. SWEBOK highlights concerns such as event-driven input models, platform-specific interface rendering, and GUI behavior modeling, issues particularly relevant to mobile GUI testing. These points help us outline core testing activities and lifecycle stages that a comprehensive taxonomy should cover, including test generation, execution, and maintenance, as well as supporting infrastructure like frameworks and tools. However, because SWEBOK provides only high-level abstraction and does not offer a detailed classification specifically for mobile GUI testing, we further refine our taxonomy using a bottom-up process. This involves a comprehensive review of our collected primary studies and several relevant secondary surveys. We systematically analyze recurring research topics, technical challenges, and contribution types to identify common themes. This dual approach allows us to construct a taxonomy that is both theoretically grounded and empirically informed, capturing the evolution of the field while maintaining alignment with established software engineering frameworks.
@@
GUI Test Record & Replay (Migration) (R&R): This topic targets at maximizing the value of manually written test scripts, which enables the automated execution of tests on different running environments without manual adaptions. Such test scripts contain the domain knowledge of human testers and can better test the critical functionalities.